Find the moment of inertia of a solid cylinder of mass 3.0kg and radius of .50m, which s free to rotate about an axis through its center
1
Expert's answer
2018-01-25T14:02:08-0500
The moment of inertia of a solid cylinder
I=mR^2/2=3×0.5^2/2=0.375 kg×m^2